# Following the Clues: Experiments on Person Re-ID using Cross-Modal Intelligence

### v1

@@ I Introduction
The proliferation of Open Data has significantly accelerated progress in artificial intelligence by providing diverse and representative datasets. However, the increasing openness of such data, combined with advancements in camera resolution, raises significant privacy concerns, particularly with visual recordings on persons in public spaces such as streets or parking areas. In particular, datasets may expose Personally Identifiable Information (PII), introducing ethical and legal risks related to person Re-Identification (Re-ID).
@@ II Related Works
Re-ID involves identifying the same individual across multiple non-interfering cameras and locations by using features like biometrics, pose, clothing, appearance, and body shape to match their identity in different frames. Recent advancements in deep learning have improved Re-ID performance, with state-of-the-art methods increasingly focusing on cross-modal representation and distance metric learning [2, 3, 8, 9] while handling prevalent challenges such as occlusions [10, 11, 12, 13] or cloth-changing, c.f. [14].
@@ III Methodology
To enhance preprocessing robustness, we automatically repair malformed LVLM outputs. To address parsing errors, we first try to parse the JSON. If a JSONDecodeError occurs, we apply automatic corrections and use the LLM phi4 (14b) [21] to fix improperly escaped quotes, handle invalid edge sets, remove trailing commas, and correct misformatted attributes. With the LLM, we use the prompt in Figure 4.
@@
$\alpha_{uv}$ $=$ (1)
$\text{softmax}_{v}\Big{(}\mathbf{a}^{\top}\text{LeakyReLU}\Big{(}$ $\mathbf{W}_{s}\mathbf{x}_{u}+\mathbf{W}_{t}\mathbf{x}_{v}+\mathbf{W}_{\mathbf{e}}\mathbf{e}_{uv}\Big{)}\Big{)},$
@@
  • LeakyReLU: Leaky Rectified Linear Unit activation
@@
$\mathcal{L}=\mathcal{L}_{ID}+\mathcal{L}_{Triplet}+\lambda\mathcal{L}_{Center}.$ (2)
@@ IV-A Setup
CUHK03-np (detected) [28, 30] contains $14,096$ images of $1,467$ identities. Each identity is captured from $2$ cameras on the CUHK campus, with an average of $4.8$ images per camera. The new protocol divides the dataset into a training set ($767$ identities) and a testing set ($700$ identities). The exact labeled split is query ($1,400$), test ($5,328$) and train ($7,368$). We use the noisy version, where bounding boxes were detected using DPM.
@@ IV-B Results
To obtain interpretability, we test edge attribution using GAtt [7], a hyperparameter-free and deterministic explanation method specifically designed for GATs. Experiments (Figure 5, bottom) indicate an averaging effect in attribution scores (similar color = similar score). Subsequently, in our setting GAtt does not provide insights about most important nodes, however filters person related information due to edge directions. The nodes are numerated based on the order obtained from the LVLM. In future work, a possible solution might be to penalize close attention weights in the loss function and/or favor simpler losses.
Figure 5: Example Re-ID query with interpretability analysis using attribution scores from GAtt [7] on node ’0: person’. The model weights were loaded from GAT / DINOv2_vitb14 – trained on Market1501/bounding_box_train [29]. The query image (0473_c4s2_050698_00.jpg) is from the query split, while the retrievals are from the gallery set bounding_box_test. Edge attributes are omitted.

### v2

@@ I Introduction
The proliferation of Open Data has significantly accelerated progress in artificial intelligence by providing diverse and representative datasets. However, the increasing openness of such data, combined with advancements in camera resolution, raises significant privacy concerns, particularly with visual recordings on persons in public spaces such as streets or parking areas. Datasets may expose Personally Identifiable Information (PII), introducing ethical and legal risks related to person Re-Identification (Re-ID). This process involves identifying the same individual across multiple non-interfering cameras and locations by using features like biometrics, pose, clothing, appearance, and body shape to match their identity in different frames.
@@ II Related Works
Recent advancements in deep learning have improved Re-ID performance, with state-of-the-art methods increasingly focusing on cross-modal representation and distance metric learning [2, 3, 8, 9] while handling prevalent challenges such as occlusions [10, 11, 12, 13] or cloth-changing, c.f. [14].
@@ III Methodology
To robustly recover valid scene graphs from malformed JSON input, the system combines a Large Language Model (LLM) (phi4 (14b) [21])-based repair function (fix_graph) with a rule-based post-processing step (fix_malformed_json). We opted for an LLM in the fixing step due to its superior adaptability and ability to handle a broader, more complex range of malformed outputs often produced by LVLMs. While simpler rule-based systems are efficient for known patterns, they often struggle with the nuanced and inconsistent errors that can arise from advanced model outputs. The LLM’s contextual understanding enables intelligent corrections that go beyond rigid syntactic rules, effectively repairing issues such as invalid edge sets or misformatted attributes that may not strictly violate JSON syntax but still disrupt the intended structure (prompt in Figure 4). After the LLM proposes a fix, fix_malformed_json applies systematic syntax cleanups – removing extraneous markers, correcting improper escaping, and restructuring edge definitions to ensure the result is both semantically coherent and syntactically valid. Notably, the fixing was rarely necessary in practice: only $30$ out of $36,036$ graphs in Market-1501, $3$ out of $14,097$ in CUHK03-NP (detected), and $6$ out of $14,096$ in CUHK03-NP (labeled) required repair, indicating high initial annotation quality and the robustness of the overall processing pipeline.
@@
$\alpha_{uv}=\operatorname{softmax}_{v}\Bigl{(}\mathbf{a}^{\top}\operatorname{LReLU}\Bigl{(}\mathbf{W}_{s}\mathbf{x}_{u}+\mathbf{W}_{t}\mathbf{x}_{v}+\mathbf{W}_{e}\mathbf{e}_{uv}\Bigr{)}\Bigr{)}$ (1)
@@
  • $\operatorname{LReLU}(x)$: Leaky Rectified Linear Unit activation
@@
$\mathcal{L}=\mathcal{L}_{Triplet}+\lambda\mathcal{L}_{Center}+\mathcal{L}_{ID}.$ (2)
@@ IV-A Setup
CUHK03-np (detected) [28, 30] contains $14,096$ images of $1,467$ identities, each captured from two cameras on the CUHK campus – averaging $4.8$ images per camera. The new protocol divides the dataset into a training set ($767$ identities) and a testing set ($700$ identities). The precise split comprises: $1,400$ query images, $5,328$ gallery (test) images, and $7,368$ training images for the labeled set, and $1,400/5,332/7,365$ for the detected split. We use the noisy (detected) version, where bounding boxes are generated using the DPM pedestrian detector.
@@ IV-B Results
Figure 5 exemplifies the ability to extract and visualize semantically meaningful, graph-structured attributes of PII from visual data. These attributes, listed in order as obtained from the LVLM, are then transformed into explicit textual graphs (e.g., “black backpack”, blue shirt”, “short black hair”), providing an interpretable understanding of what can make a person re-identifiable beyond pose and face biometrics. Finding these graphs can enhance the safety of public datasets from a privacy perspective in several ways:
  • Complementary Risk Assesment: The explicit textual attributes within these graphs allow for a systematic and granular evaluation of Re-ID risks. This goes beyond merely detecting faces to understanding specific visual cues (like clothing details or carried objects) that an adversary could exploit. By understanding these specific semantic links, the framework can model a realistic Re-ID threat scenario that consider both visual and textual attack vectors. This capability helps data publishers assess how individuals might be identified and evaluate datasets for their robustness against sophisticated Re-ID attempts, which is crucial for compliant data release.
  • Targeted Anonymization: Once these interpretable PII attributes are identified through the graph representations, data publishers can implement more precise and effective anonymization strategies. Instead of broad blurring or pixelation that might destroy valuable data utility, specific features contributing to re-identifiability can be targeted for masking or alteration. This allows for a balance between data utility for research and individual privacy protection.
To further invastigate this interpretability, we test edge attribution using GAtt [7], a hyperparameter-free and deterministic explanation method specifically designed for GATs. Experiments (Figure 5, bottom) indicate an averaging effect in attribution scores (similar color = similar score). Subsequently, in our setting GAtt does not provide insights about most important nodes, however filters person-related information due to edge directions. In future work, a possible solution might be to penalize close attention weights in the loss function and/or favor simpler losses.
Figure 5: Example Re-ID query with interpretability analysis using attribution scores from GAtt [7] on node ’0: person’. This figure illustrates a query image and its top retrieved matches, alongside the textual graph representations (nodes and attributes) and their corresponding attribution scores. Due to GAtt’s methodology, non-contributing nodes are omitted in these graph plots. The model weights were loaded from GAT / DINOv2_vitb14 – trained on Market1501/bounding_box_train [29]. The query image (0473_c4s2_050698_00.jpg) is from the query split, and the retrievals are from the gallery set bounding_box_test. Edge attributes are omitted for better readability. The nodes are numerated based on the order obtained from the LVLM.